# Time-Aware and Transition-Semantic Graph Neural Networks for Interpretable Predictive Business Process Monitoring  [^1] The full codebase is available at https://github.com/skyocean/TemporalAwareGNNs-NextEvent

### v1

@@ 5.2 Experiment Setup
All models were implemented using PyTorch Geometric 2.5.3 (Fey and Lenssen, 2019) and executed on a machine equipped with an Intel(R) Core(TM) i9-8950HK CPU (6 cores, 12 threads) and 32 GB of RAM. No GPU acceleration was employed.
@@ 5.2.1 Global Configuration
For GAT-based models, we adopt a deeper architecture to accommodate the increased model capacity. Node attributes are processed using a GAT layer with 32 hidden units, while event labels are embedded into 64 dimensional vectors and passed through a GAT layer of hidden size 128. The fusion layer uses 256 hidden units, with 4 attention heads applied consistently across all GAT layers. Transition aware models embed edge types into a 32 dimensional vector. A fixed learning rate of 0.001 and Adam optimizer are used. For time decay models, the decay coefficient $\lambda$ is selected from {0.01, 0.1, 0.3, 0.7}, dynamically adjusted based on the median length of event sequences in each dataset.
@@ 5.2.2 Evaluation Metrics
  • Accuracy (Acc): Measures the proportion of test instances where the top-1 predicted label exactly matches the ground truth.
@@ 6.1 Prefix-Based GCN Performance Analysis
The performance of our next event prediction model was evaluated across five benchmark datasets using top-k accuracy metrics (k = 1, 3, 5), under varying prefix lengths ranging from 2 to 20 events. Figure. 1 and Table. 2 present a comprehensive overview of the results, highlighting consistent trends across datasets.
@@
Overall, most datasets achieved peak prediction accuracy with relatively short prefixes (typically 3-5 events), regardless of the underlying sequence length distributions. This indicates that the next event is most influenced by the most recent few steps in the trace, and that longer histories may introduce diminishing returns or even noise. For instance, in BPI12W, performance dropped from 0.8531 at prefix 5 to 0.7843 and 0.7823 at prefixes 10 and 20, respectively. A similar trend was observed across other datasets. This suggests that short range dependencies dominate in most real world cases, so incorporating more distant context may dilute the signal or overfit to idiosyncratic cases.
This trend may stem from two interacting factors. First, as prefix length increases, the number of available samples decreases due to fewer long traces, introducing sample bias and instability. Especially, the fixed-size output space remains the same and large(e.g., 36 for BPI12, 19 for BPI12W) Second, the GCN model lacks an attention mechanism, treating all input nodes equally. As a result, important past events may be undervalued, while noisy or less relevant events may be overemphasized.
We also observe substantial gains from Top-1 to Top-3 accuracy, particularly in datasets with more complex process structures. The average improvements from Top-1 to Top-3 were 0.151 (BPI12W), 0.113 (BPI12), 0.201 (BPI13i), 0.244 (BPI13c), and 0.038 (Helpdesk). This reflects greater ambiguity in next event prediction under complex workflows, where the top prediction may miss but the correct label still appears in a small candidate set. Notably, Top-5 accuracy remains high across all prefix lengths and datasets (e.g., BPI12W: 0.9995 at prefix 3, 0.9937 at prefix 20), supporting the model’s utility for decision support even in low confidence contexts.
@@ 6.2 GAT Variants Performance Analysis
Table 3 presents the performance of four GAT variants evaluated across five datasets using Top-1/3/5 accuracy and Damerau–Levenshtein (DL) distance. The best scores are highlighted for each metric.
Table 3: Top-k accuracy and Damerau–Levenshtein distance (DL) of GAT variants.
@@
Model Top-1 Top-3 Top-5 DL Top-1 Top-3 Top-5 DL
GAT-T 0.6847 0.9121 0.9706 0.6986 0.6188 0.8665 0.9793 0.6657
GAT-TD 0.6989 0.9049 0.9673 0.6895 0.4932 0.8044 0.9615 0.5069
GAT-$T$ 0.6922 0.9113 0.9721 0.6954 0.6360 0.8637 0.9686 0.6521
GAT-TDTE 0.6995 0.8904 0.9577 0.6858 0.4797 0.7852 0.9464 0.5004
Helpdesk
Model Top-1 Top-3 Top-5 DL
GAT-T 0.9570 0.9915 0.9967 0.9622
GAT-TD 0.7531 0.9509 0.9911 0.7395
GAT-$T$ 0.9584 0.9920 0.9955 0.9703
GAT-TDTE 0.7516 0.9457 0.9880 0.7371
  • For GAT-TD and GAT-TDTE, the best-performing $\lambda$ was selected per dataset from 0.01,0.1,0.3,0.7{0.01,0.1,0.3,0.7}. Optimal values were: BPI12 and BPI12W — 0.7; BPI13i — 0.3; BPI13c — 0.1; Helpdesk — 0.01.
6.2.1 GAT with Edge-Aware Attention Performance Analysis
Compared to the prefix based GCN models, both baseline GAT variants (GAT-T and GAT-TT) generally achieve higher Top-1 accuracy, with the only exception being GAT-T on BPI13c. This improvement indicates that the attention mechanism in GAT models effectively identifies and prioritizes historically relevant events beyond the limited prefix window. By considering the full event graph, the model is better positioned to capture long range dependencies that static GCNs may fail to exploit.
Across all datasets, models that incorporate transition type embeddings in their edge features (GAT-TT) show accuracy improvements over their time only counterparts. This supports the hypothesis that capturing semantic transitions between events enhances the model’s ability to distinguish among plausible next steps, especially in complex or noisy process flows.
@@ 6.2.1 GAT with Edge-Aware Attention Performance Analysis
The only notable performance drop occurs on BPI13c, where GAT-T underperforms the GCN baseline. This dataset has a short median sequence length of three, a wide temporal span, and low event label diversity. Under such conditions, full graph attention based solely on local time differences may offer limited utility. However, when semantic edge features are introduced, as in GAT-TT, performance improves. This suggests that effective attention in sparse and temporally stretched logs requires richer contextual information beyond temporal proximity.
@@ 6.2.2 GAT with Dynamic Temporal Windows Performance Analysis
For long sequence datasets such as BPI12 and BPI12W, we observed a modest drop in accuracy (1.4% to 4%) when using decay based attention mechanisms (GAT-TD, GAT-TDTE). Although this reduction may seem counterintuitive, further analysis revealed that tuning the decay strength $\lambda$ mitigates this effect. Specifically, we found that higher decay values (e.g., $\lambda=0.7$) yielded better results on these datasets, where long traces often include irrelevant historical noise. Rather than indicating a flaw in the mechanism, this emphasizes the importance of calibrating decay to match the temporal structure of each dataset.
@@
This interpretation is further supported by the sharp accuracy drop observed in Helpdesk and BPI13c. Both datasets contain very short sequences (median length three), where most historical events are only one or two steps away. However, despite similar sequence lengths, Helpdesk has minimal time gaps between events, whereas BPI13c features larger global time differences. Consequently, a low decay strength ($\lambda=0.01$) works best for Helpdesk, while BPI13c benefits from a slightly higher value ($\lambda=0.1$). In both cases, decay must be delicately tuned: if too strong, it suppresses even nearby useful signals. These observations confirm that the effectiveness of temporal decay hinges on both sequence depth and absolute time dispersion.
Finally, we observe a meaningful interaction between temporal decay and transition type embeddings. In the mid length dataset BPI13i, incorporating transition type edge embeddings alongside temporal decay (GAT-TDTE) results in higher accuracy compared to using decay alone (GAT-TD). In contrast, this combination leads to performance degradation in both shorter and longer sequences. These results suggest that transition based edge features may intensify the temporal attenuation effect, which is beneficial when sequence depth and temporal dispersion are well balanced, but may suppress relevant information in regimes with limited or excessive context.
@@ 6.3.1 Dynamic Temporal Windows
Figure 2(a) visualizes how node importance, computed by combining attention weights with the time decay factor, varies across node positions (x axis) and sequence lengths (y axis). For interpretability, we divide sequences into discrete length bins of five events, ranging from 0 to 4 up to 90 to 94. The attention consistently concentrates on the final five to ten events, forming prediction focused windows regardless of total sequence length. For shorter sequences (10 to 29 events), attention is moderately distributed across the mid to late regions. In contrast, longer sequences (over 55 events) display sharp, localized focus on the last two to three events, with peak importance scores exceeding 0.6. This spatial shift confirms that the time decay mechanism constructs prediction optimized attention windows by amplifying recent events while attenuating distant historical context.
@@ 6.4 Comparative Discussion
We compare our model’s accuracy and Damerau–Levenshtein (DL) distance scores with prior approaches across five benchmark datasets, as reported in Table 5 and Table 5. Our framework delivers competitive results, with particularly strong performance on the Helpdesk dataset. While TACO (Rama-Maneiro et al., 2023) reports higher accuracy on structurally complex datasets such as BPI12 and BPI13, several methodological differences account for this gap and also highlight the strengths of our approach.
@@
  • Pas. = Pasquadibisceglie et al. (2019), Tax = Tax et al. (2017), Cam. = Camargo et al. (2019),Khan = Khan et al. (2018), Ev. = Evermann et al. (2017), Mau. = Di Mauro et al. (2019), Theis = Theis and Darabi (2019), TACO = Rama-Maneiro et al. (2023), Fran. =Di Francescomarino et al. (2017)
7 Conclusion
@@
Our models retain the full event history and jointly predict the next event and its substatus, enabling a more expressive and process aware output space compared to previous approaches. Empirical evaluations across five public datasets show that global attention through GAT improves Top-1 accuracy by an average of 5.5 percent over prefix based GCNs. The addition of time decay attention provides a further gain of 1.5 percent in sequences of medium length with high temporal variability. Edge embeddings informed by transition semantics enhance the model’s ability to distinguish structurally similar but semantically distinct traces. High Top-3 and Top-5 accuracies (exceeding 86%), along with superior Damerau–Levenshtein alignment, confirm both event level prediction accuracy and sequence level coherence.
@@
To support transparency and actionable decision making in predictive business process monitoring, we introduce a suite of interpretability visualizations that explain how dynamic temporal attention windows and transition semantics influence predictions across sequences of varying length. Quantitative analysis shows a consistent contraction of attention span from 34 to 20 percent as sequence length increases, indicating that the model effectively suppresses irrelevant context while maintaining predictive accuracy. Edge attention correlation analysis ($r=0.74$) confirms that transition informed edge embeddings meaningfully influence attention, particularly in sequences with structural ambiguity. These interpretability tools support practical use cases such as bottleneck detection and compliance auditing, while visual decompositions presented in  Appendix A offer domain experts a means to investigate prediction behavior on individual cases. Taken together, these contributions advance the state of predictive process monitoring by combining accuracy, interpretability, and practical deployability in a unified graph-based framework.
Finally, our approach is generalizable beyond business process monitoring to any domain involving sequential, timestamped event data, such as healthcare workflows or IT service logs. Future work may explore hierarchical attention mechanisms for nested process structures, AutoML driven hypermodels for dataset specific adaptation, and alternative forms of temporal awareness to better address highly irregular sequences. In addition, combining our semantic edge embeddings with TACO’s state accumulation strategy, such as incorporating token counts as node features, may offer complementary strengths. Integrating these advancements into practitioner oriented tools could further bridge predictive analytics with operational process optimization.

### v2

@@ 5.2 Experiment Setup
All models were implemented using PyTorch Geometric 2.5.3 (FeyLenssen2019) and executed on a machine equipped with an Intel(R) Core(TM) i9-8950HK CPU (6 cores, 12 threads) and 32 GB of RAM. No GPU acceleration was employed.
@@ 5.2.1 Global Configuration
For GAT-based models, we adopt a deeper architecture to accommodate the increased model capacity. Node attributes are processed using a GAT layer with 32 hidden units, while event labels are embedded into 64 dimensional vectors and passed through a GAT layer of hidden size 128. The fusion layer uses 256 hidden units, with 4 attention heads applied consistently across all GAT layers. Transition aware models embed edge types into a 32 dimensional vector. A fixed learning rate of 0.001 and Adam optimizer are used.
5.2.2 Decay Coefficient ($\lambda$) Analysis
For time-decay models, we evaluate $\lambda$ over the fixed candidate set $\{0.01,0.1,0.3,0.7\}$ for each dataset. This strategy reflects the temporal diversity of event logs, where case lengths and attribute ranges vary substantially. In the reported results, we present the best-performing configuration of $\lambda$ for each dataset. This procedure not only ensures fair comparison across datasets but also allows us to assess the sensitivity of performance to the decay factor, thereby clarifying the role of $\lambda$ under different temporal profiles.
@@ 5.2.3 Evaluation Metrics
  • Accuracy (Acc): Measures the proportion of test instances where the top-1 predicted label exactly matches the ground truth. For completeness, and to account for class imbalance across logs, we also report precision (P), recall (R), and weighted F1-score (F1) for GAT models (see Table 3), though our primary focus remains on accuracy and sequence-level alignment metrics.
@@ 6.1 Prefix-Based GCN Performance Analysis
The performance of our next event prediction model was evaluated across six benchmark datasets using top-k accuracy metrics (k = 1, 3, 5), under varying prefix lengths ranging from 2 to 20 events. Figure. 3 and Table. 2 present a comprehensive overview of the results, highlighting consistent trends across datasets.
@@
Overall, most datasets achieved peak prediction accuracy with relatively short prefixes (typically 3-5 events), regardless of the underlying sequence length distributions. This indicates that the next event is most influenced by the most recent few steps in the trace, and that longer histories may introduce diminishing returns or even noise. For instance, in BPI12W, performance dropped from 0.8531 at prefix 5 to 0.7843 and 0.7823 at prefixes 10 and 20, respectively. A similar trend was observed across other datasets, reflecting two interacting factors. First, as prefix length increases, the number of available samples decreases due to fewer long traces, introducing sample bias and instability. Especially, the fixed-size output space remains the same and large(e.g., 36 for BPI12, 19 for BPI12W) Second, the GCN model lacks an attention mechanism, treating all input nodes equally. As a result, important past events may be undervalued, while noisy or less relevant events may be overemphasized.
BPI20 presents an exception: accuracy remains high at longer prefixes (e.g., 0.9732 at prefix 10), rather than dropping. This pattern reflects its structural profile—short maximum length (21 events), but relatively large median length (8 events), meaning that most traces are already moderately long. In addition, BPI20 contains 29 prediction classes and richer sequence-level attributes but minimal event-level features. These factors together make longer prefixes informative rather than noisy, since extending the input better exposes the case-level context needed to discriminate among many classes. Thus, BPI20 complements the other benchmarks by highlighting conditions where longer-range prefixes strengthen rather than weaken predictive performance.
We also observe substantial gains from Top-1 to Top-3 accuracy, particularly in datasets with more complex process structures. The average improvements from Top-1 to Top-3 were 0.151 (BPI12W), 0.113 (BPI12), 0.201 (BPI13i), 0.244 (BPI13c), 0.038 (Helpdesk), 0.058 (BPI20). This reflects greater ambiguity in next event prediction under complex workflows, where the top prediction may miss but the correct label still appears in a small candidate set. Notably, Top-5 accuracy remains high across all prefix lengths and datasets (e.g., BPI12W: 0.9995 at prefix 3, 0.9937 at prefix 20), supporting the model’s utility for decision support even in low confidence contexts.
@@ 6.2 GAT Variants Performance Analysis
Table 3 reports the performance of four GAT variants evaluated across six datasets. Metrics include Top-1/3/5 accuracy, precision, recall, weighted F1, and Damerau–Levenshtein (DL) distance. Best-performing values are highlighted for accuracy and DL measures.
Table 3: Top-k accuracy, Damerau–Levenshtein distance (DL), Precision(P), Recall(R) and weighted F1 (F1) of GAT Model(M) variants.
@@
M Top-1 Top-3 Top-5 DL P R F1 Top-1 Top-3 Top-5 DL P R F1
T 0.6847 0.9121 0.9706 0.6986 0.673 0.6847 0.6752 0.6188 0.8665 0.9793 0.6657 0.624 0.621 0.6204
TD 0.6989 0.9049 0.9673 0.6895 0.6787 0.693 0.6807 0.4932 0.8044 0.9615 0.5069 0.552 0.5032 0.489
$T$ 0.6922 0.9113 0.9721 0.6954 0.6847 0.6894 0.6916 0.6360 0.8637 0.9686 0.6521 0.5988 0.601 0.6265
TDTE 0.6995 0.8904 0.9577 0.6858 0.6882 0.6842 0.6958 0.4797 0.7852 0.9464 0.5004 0.5072 0.4912 0.4705
Helpdesk BPI20
M Top-1 Top-3 Top-5 DL P R F1 Top-1 -Top-3 Top-5 DL P R F1
T 0.9570 0.9915 0.9967 0.9622 0.9539 0.9495 0.9482 0.8295 0.9568 0.9903 0.8184 0.806 0.8143 0.804
TD 0.7531 0.9509 0.9911 0.7395 0.7782 0.7412 0.7498 0.7529 0.927 0.9791 0.732 0.8027 0.7211 0.7458
$T$ 0.9584 0.9920 0.9955 0.9703 0.9586 0.9584 0.9582 0.8439 0.9591 0.9953 0.843 0.8234 0.8204 0.8119
TDTE 0.7516 0.9457 0.9880 0.7371 0.7899 0.736 0.7508 0.7515 0.9056 0.9627 0.724 0.8024 0.7334 0.7394
  • For GAT-TD and GAT-TDTE, we evaluate the decay coefficient at four preset values, $\lambda\in\{0.01,\,0.1,\,0.3,\,0.7\}$, on each dataset and report the best-performing value. Selected $\lambda$ per dataset: BPI12/BPI12W: 0.7; BPI13i: 0.3; BPI13c: 0.1; Helpdesk/BPI20: 0.01.The same four-value evaluation is applied uniformly across datasets to assess sensitivity; no bespoke tuning is performed.
6.2.1 Comparative Performance of GATs and Prefix-GCNs across Datasets
Across five of the six datasets, at least one GAT variant outperforms prefix-GCNs in Top-1 accuracy. This improvement indicates that the attention mechanism in GAT models effectively identifies and prioritizes historically relevant events beyond the limited prefix window. By considering the full event graph, the model can capture long-range dependencies that prefix-based GCNs, limited to partial traces, may fail to exploit. The only exception is BPI20. BPI20 combines many next-event classes (29) with a relative small sample (2,099 traces), moderate sequence length (median 8; max 21), and minimal event-level features (versus several sequence-level attributes). In this regime, per-edge signals are thin: local time gaps and transition types carry limited discriminative power across 29 classes, so full-graph attention tends to diffuse evidence and add variance in a low-sample, high-cardinality setting. Dynamic temporal decay can also de-emphasize early approvals that remain decisive later in the case. By contrast, a long-prefix GCN acts as a low-variance aggregator: once the prefix covers most of the trace (e.g., P=10P{=}10 is near full length for BPI20), it captures the decisive late-stage context without relying on per-edge attention. This explains why short-prefix GCNs and GATs underperform on BPI20, while near-full-length GCN achieves the best Top-1 (Table 2, Table 3). In summary, global attention helps when event-level signals are sufficiently informative to benefit from selective weighting. When class cardinality is high, samples are limited, and per-event features are sparse (as in BPI20), long-prefix GCNs remain a strong, pragmatic choice.
6.2.2 GAT with Edge-Aware Attention: Performance Analysis
Across all datasets, adding transition-type embeddings to edges (GAT-TT) consistently improves over time-only attention (GAT-T) in both Top-1 accuracy and weighted F1. The gains are modest but uniform (0.11.60.1–1.6 percentage points in accuracy across logs), confirming that semantic transitions help discriminate among plausible next steps. Weighted F1 closely tracks Top-1 trends, showing that these improvements extend beyond majority-class effects.
@@ 6.2.2 GAT with Edge-Aware Attention: Performance Analysis
BPI13c illustrates this effect most clearly: while plain GAT-T lags behind prefix GCNs, GAT-TT recovers and attains the best GAT Top-1 and F1 on the log. This dataset has a short median sequence length of three, a wide temporal span, and low event label diversity. Under such conditions, full graph attention based solely on local time differences may offer limited utility. However, when semantic edge features are introduced, as in GAT-TT, performance improves. This suggests that effective attention in sparse and temporally stretched logs requires richer contextual information beyond temporal proximity.
BPI20 is the exception where prefix-GCNs outperform all GAT variants (see Section 6.2.1). Still, within the GAT family, semantics remain beneficial: GAT-TT surpasses GAT-T in both Top-1 (0.844 vs. 0.830) and weighted F1 (0.812 vs. 0.804). This indicates that while global attention is not the best paradigm under high-class cardinality, low sample size, and minimal event-level features, semantic transitions continue to strengthen per-edge signals.
@@ 6.2.3 GAT with Dynamic Temporal Windows Performance Analysis
For long sequence datasets such as BPI12 and BPI12W, we observed a modest drop in accuracy (1.4% to 4%) when using decay based attention mechanisms (GAT-TD, GAT-TDTE). Although this reduction may seem counterintuitive, further analysis revealed that selecting the decay strength $\lambda$ mitigates this effect. Specifically, we found that higher decay values (e.g., $\lambda=0.7$) yielded better results on these datasets, where long traces often include irrelevant historical noise. Rather than indicating a flaw in the mechanism, this emphasizes the importance of calibrating decay to match the temporal structure of each dataset.
@@
This interpretation is further supported by the sharp accuracy drop observed in Helpdesk and BPI13c. Both datasets contain very short sequences (median length three), where most historical events are only one or two steps away. However, despite similar sequence lengths, Helpdesk has minimal time gaps between events, whereas BPI13c features larger global time differences. Consequently, a low decay strength ($\lambda=0.01$) works best for Helpdesk, while BPI13c benefits from a slightly higher value ($\lambda=0.1$). In both cases, decay must be delicately selected: if too strong, it suppresses even nearby useful signals. These observations confirm that the effectiveness of temporal decay hinges on both sequence depth and absolute time dispersion.
We observe a meaningful interaction between temporal decay and transition type embeddings. In the mid length dataset BPI13i, incorporating transition type edge embeddings alongside temporal decay (GAT-TDTE) results in higher accuracy compared to using decay alone (GAT-TD). In contrast, this combination leads to performance degradation in both shorter and longer sequences. These results suggest that transition based edge features may intensify the temporal attenuation effect, which is beneficial when sequence depth and temporal dispersion are well balanced, but may suppress relevant information in regimes with limited or excessive context.
Finally, BPI20 shows limited benefit from temporal decay: with short traces, high class cardinality, and sparse event-level features, decay can further dilute already weak per-edge signals. This explains why GAT-TD and GAT-TDTE underperform their non-decay counterparts, even though $\lambda=0.01$ minimizes the loss.
Overall, these results confirm that while the exact value of $\lambda$ can shift across datasets, model performance is robust within the tested range $\{0.01,0.1,0.3,0.7\}$, and the observed trends validate the role of temporal decay as a consistent mechanism rather than a dataset-specific tuning artifact.
@@ 6.3.1 Dynamic Temporal Windows
Figure 4(a) visualizes how node importance, computed by combining attention weights with the time decay factor, varies across node positions (x axis) and sequence lengths (y axis). For interpretability, we divide sequences into discrete length bins of six events, ranging from 0 to 4 up to 90 to 94. The attention consistently concentrates on the final six to ten events, forming prediction focused windows regardless of total sequence length. For shorter sequences (10 to 29 events), attention is moderately distributed across the mid to late regions. In contrast, longer sequences (over 55 events) display sharp, localized focus on the last two to three events, with peak importance scores exceeding 0.6. This spatial shift confirms that the time decay mechanism constructs prediction optimized attention windows by amplifying recent events while attenuating distant historical context.
@@ 6.3.2 Edge Attention Correlation (Global Sensitivity)
Together, these results strengthen the interpretability argument: by linking attention weights to explicit transition categories, the analysis reveals which process relations consistently drive prediction outcomes. This makes the attention mechanism not only a performance enhancer but also an interpretable lens, exposing how temporal and semantic dependencies interact to shape model focus across cases.
@@ 6.4 Comparative Discussion
We compare our model’s accuracy and Damerau–Levenshtein (DL) distance scores with prior approaches across six benchmark datasets, as reported in Table 5 and Table 5. Our framework delivers competitive results, with particularly strong performance on the Helpdesk dataset. While TACO (rama2023embedding) reports higher accuracy on structurally complex datasets such as BPI12 and BPI13, several methodological differences account for this gap and also highlight the strengths of our approach.
@@
BPI20 - - - - -
  • Pas. = pasquadibisceglie2019using, Tax = tax2017predictive, Cam. = camargo2019learning,Khan = khan2018memory, Ev. = evermann2017predicting, Mau. = di2019activity, Theis = theis2019decay, TACO = rama2023embedding, Fran. =di2017eye
7 Conclusion and Discussion
@@
Our models retain the full event history and jointly predict the next event and its substatus, enabling a more expressive and process aware output space compared to previous approaches. Empirical evaluations across six public datasets show that global attention through GAT improves Top-1 accuracy by an average of 5.5 percent over prefix based GCNs. The addition of time decay attention provides a further gain of 1.5 percent in sequences of medium length with high temporal variability. Edge embeddings informed by transition semantics enhance the model’s ability to distinguish structurally similar but semantically distinct traces. High Top-3 and Top-5 accuracies (exceeding 86%), along with superior Damerau–Levenshtein alignment, confirm both event level prediction accuracy and sequence level coherence. Crucially, we also delineate the boundary of global attention: in a regime with high class cardinality, sparse event-level features, and short-to-moderate traces (BPI20), long-prefix GCNs outperform GATs. The unified framework and its interpretability decomposition make the regime explicit and provide prescriptive guidance: use GATs for longer, feature-rich traces where selective weighting pays off; use long-prefix GCNs when per-event signal is weak and the class space is large. The contribution is therefore not only improved averages, but a clear, data-driven rule for model selection across process regimes.
@@
To support transparency and actionable decision making in PBPM, we introduce a suite of interpretability visualizations that explain how dynamic temporal attention windows and transition semantics influence predictions across sequences of varying length. Quantitative analysis shows a consistent contraction of attention span from 34 to 20 percent as sequence length increases, indicating that the model effectively suppresses irrelevant context while maintaining predictive accuracy. Edge attention correlation analysis ($r=0.74$) confirms that transition informed edge embeddings meaningfully influence attention, particularly in sequences with structural ambiguity. These interpretability tools support practical use cases such as bottleneck detection and compliance auditing, while visual decompositions presented in  Appendix A offer domain experts a means to investigate prediction behavior on individual cases. Taken together, these contributions advance the state of predictive process monitoring by combining accuracy, interpretability, and practical deployability in a unified graph-based framework.
While the proposed framework advances PBPM through a unified design that integrates prefix-based GCNs and full-trace GATs with temporal decay and transition semantics, several limitations remain and point to future work. First, we deliberately fixed most hyperparameters globally to ensure comparability across models and datasets, but more extensive hyperparameter optimization could further improve performance. Exploring such optimization in conjunction with automated search techniques (e.g., AutoML) would help clarify which architectural components benefit most from dataset-specific tuning. Second, in this study, categorical features were encoded using one-hot representations across all models to ensure a fair and controlled comparison. This choice limits expressiveness but avoids introducing confounding variation in our architectural benchmarks. In future work, alternative encoding strategies (e.g., embeddings or frequency-based encodings) could be explored to examine which model architectures benefit most from richer feature representations. Third, while our evaluation relies on bucket-based prefix sampling to balance short, medium, and long prefixes, an alternative is to enumerate all prefixes and focus on last-event prediction. We consider this complementary setup an interesting extension for future work, particularly for scenarios where complete trace coverage is prioritized over prefix diversity. Four, one limitation of our approach is the assumption of complete and consistent transition-type labeling. While in benchmark event logs transition categories are derived deterministically from observed event pairs, in real-world applications incomplete or noisy labeling may reduce the reliability of transition embeddings. Extending the framework to handle uncertain or probabilistic transition types is an important direction for future work. Fifth, we evaluated on six benchmark datasets spanning multiple domains, but future studies could expand coverage to additional real-world logs to further stress-test generalizability. Moreover, our current models operate at the sequence level; extending them to hierarchical or multi-instance settings (e.g., nested or interacting processes) could broaden applicability. In addition, combining our semantic edge embeddings with TACO’s state accumulation strategy, such as incorporating token counts as node features, may offer complementary strengths. Integrating these advancements into practitioner oriented tools could further bridge predictive analytics with operational process optimization.